OpenStudy (anonymous):

Geometric means of 6 and 12? I know the answer is either: 36, 72, 2 radical 6, or 6 radical 2. I just need someone to show me how to solve this.. Geometry isn't explaining it enough for me, and I don't remember how to solve square roots from Algebra. :/ So some help would be nice.. thanks. c:

jimthompson5910 (jim_thompson5910):

to find the geometric mean of two numbers, multiply them and then take the square root of that result

jimthompson5910 (jim_thompson5910):

remember to simplify fully

OpenStudy (anonymous):

I know that I have to multiply them, but I don't understand how to get the square root. :/

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

How do I find the square root of 72?

jimthompson5910 (jim_thompson5910):

you simplify that, you don't need to evaluate it

OpenStudy (anonymous):

I don't understand what you're trying to say. D:

jimthompson5910 (jim_thompson5910):

sqrt(72) = sqrt(36*2) sqrt(72) = sqrt(36)*sqrt(2) sqrt(72) = ???

OpenStudy (anonymous):

To get the geometric means of 6 and 12, I know I need to do \[\sqrt{ab}\] but I don't understand what to do AFTER that.

OpenStudy (anonymous):

oh. So the square root of 36 is 6.. so it would be \[6\sqrt{2}\] right?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Okay, thanks. :D

jimthompson5910 (jim_thompson5910):

np